OpenStudy (anonymous):

I urgently need help or I WILL fail my class :'( please explain this to me.. Solve each equation for y. 3. 4x + 2y = 6 4. -3x – y = 4

OpenStudy (dumbcow):

haha please don;t be so dramatic goal is to isolate y -move the x term to other side - divide by the coefficient ex: 7x -3y = 9 -7x -7x --> -3y = 9-7x --> y = (9-7x)/-3 = -3 +(7/3)x

OpenStudy (kymber):

You just get y alone by undoing operations First you subtract the 4x Then you divide two to make y alone.